OpenStudy (anonymous):

how do you simplify (x^2 + 4x – 3)(2x^2 + x + 6)?

OpenStudy (neo92):

u mean factorize?

OpenStudy (anonymous):

the directions just say to simplify and put in standard form

OpenStudy (amistre64):

id say FOIL, but that would be pointless eh

jimthompson5910 (jim_thompson5910):

hint: (x^2 + 4x – 3)(2x^2 + x + 6) x^2(2x^2 + x + 6) + 4x(2x^2 + x + 6) -3(2x^2 + x + 6) From here, distribute and simplify

OpenStudy (anonymous):

well I'm working with multiplying polynomials if that's any help...

OpenStudy (amistre64):

jims got a good setup going

OpenStudy (anonymous):

is the first part 2x^4 + x^3 + 6x^2

jimthompson5910 (jim_thompson5910):

you got it

jimthompson5910 (jim_thompson5910):

what about the second piece?

OpenStudy (anonymous):

8x^3 + 4x^2 + 24x? not too sure on that one

jimthompson5910 (jim_thompson5910):

right again

jimthompson5910 (jim_thompson5910):

now the last part

OpenStudy (anonymous):

-6x^2 - 3x -9?

jimthompson5910 (jim_thompson5910):

close

OpenStudy (anonymous):

sorry if I'm taking to slow haha! & let me retry

jimthompson5910 (jim_thompson5910):

-3 times -6 is not -9 though

jimthompson5910 (jim_thompson5910):

you're doing just fine, no need to hurry

OpenStudy (anonymous):

-12?

jimthompson5910 (jim_thompson5910):

3 times 6 = ???

OpenStudy (anonymous):

oh wait -18?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

does a negative times a negative equal a positive?

jimthompson5910 (jim_thompson5910):

oh sry, i meant -3 times 6

jimthompson5910 (jim_thompson5910):

not sure why I wrote -6, it's actually positive

jimthompson5910 (jim_thompson5910):

Anyways, so we have these steps so far (x^2 + 4x – 3)(2x^2 + x + 6) x^2(2x^2 + x + 6) + 4x(2x^2 + x + 6) -3(2x^2 + x + 6) 2x^4+x^3+6x^2 + 8x^3+4x^2+24x -6x^2-3x-18

jimthompson5910 (jim_thompson5910):

the last step is to combine like terms

OpenStudy (anonymous):

6x^10 + 5x^5 +21x 18?

OpenStudy (anonymous):

21x - 18*

jimthompson5910 (jim_thompson5910):

no, you only add terms that have the same exponent

jimthompson5910 (jim_thompson5910):

ex: you can only add x^3 to 8x^3 to get x^3 + 8x^3 = 1x^3 + 8x^3 = (1+8)x^3 = 9x^3 So x^3 + 8x^3 = 9x^3

jimthompson5910 (jim_thompson5910):

and adding these terms does NOT change the exponent

OpenStudy (anonymous):

oh yeah! that completly slipped my mind! im so sorry

jimthompson5910 (jim_thompson5910):

no worries

jimthompson5910 (jim_thompson5910):

after you correct, tell me what you get

OpenStudy (anonymous):

4x^2 +2x^4 + 9x^3 +24x + 3x -18?

jimthompson5910 (jim_thompson5910):

how did you get 24x?

jimthompson5910 (jim_thompson5910):

oh wait, i see now, it should be 24x - 3x Combine that and tell me what you get